# The Primacy of Magnitude in Low-Rank Adaptation

### v1

@@ Abstract
Low-Rank Adaptation (LoRA) offers a parameter-efficient paradigm for tuning large models. While recent spectral initialization methods improve convergence and performance over the naive “Noise & Zeros” scheme, their extra computational and storage overhead undermines efficiency. In this paper, we establish update magnitude as the fundamental driver of LoRA performance and propose LoRAM [^1] Code will be released at this page, a magnitude-driven “Basis & Basis” initialization scheme that matches spectral methods without their inefficiencies. Our key contributions are threefold: (i) Magnitude of weight updates determines convergence. We prove low-rank structures intrinsically bound update magnitudes, unifying hyperparameter tuning in learning rate, scaling factor, and initialization as mechanisms to optimize magnitude regulation. (ii) Spectral initialization succeeds via magnitude amplification. We demystify that the presumed knowledge-driven benefit of the spectral component essentially arises from the boost in the weight update magnitude. (iii) A novel and compact initialization strategy, LoRAM, scales deterministic orthogonal bases using pretrained weight magnitudes to simulate spectral gains. Extensive experiments show that LoRAM serves as a strong baseline, retaining the full efficiency of LoRA while matching or outperforming spectral initialization across benchmarks.
@@ 1 Introduction
The rise of large pretrained models Brown2020LanguageMA ; OpenAI2023GPT4TR ; touvron2023llama ; bai2023qwen ; guo2025deepseek has driven urgent needs for parameter-efficient fine-tuning (PEFT) methods peft ; Hu2021LoRALA ; Lester2021ThePO ; He2021TowardsAU ; Edalati2022KronAPE ; zhang2025parameter . Among these, Low-Rank Adaptation (LoRA) Hu2021LoRALA stands out for its efficiency, flexibility, and stability. By freezing pretrained weights and injecting trainable low-rank matrices, LoRA enables the update of less than 1% of the parameters, significantly reducing memory and compute costs. Its plug-and-play nature achieves easy integration into diverse models, facilitating model sharing and federated learning SLoRA ; FeDeRA . Furthermore, LoRA helps prevent catastrophic forgetting biderman2024lora , making it suitable for continuous learning. These advantages have led to its wide adoption in NLP liu2022few ; Ding2022DeltaTA ; sun2023comparative ; zhao2024lora and multimodal applications ye2023mplug ; guo2023animatediff ; blattmann2023stable ; ruiz2023dreambooth ; zhang2023towards .
@@ 2.1 Preliminaries and Notations
where $B\in\mathbb{R}^{n\times r}$, $A\in\mathbb{R}^{r\times m}$ are trainable low-rank matrices with $r\ll\min(n,m)$, and $\alpha$ scales the update magnitude. The initialization term $B^{(0)}A^{(0)}$ can be absorbed into $W$ for computation convenience, as illustrated in Figure 1(b). Given a loss function $L$, LoRA updates are computed as:
@@ 2.2 Effect of Hyperparameters on Update Magnitude
Initialization $A^{(0)}=A_{\text{init}},B^{(0)}=B_{\text{init}}\end{aligned}$ $\tilde{A}^{(0)}=\alpha_{A}A_{\text{init}},\tilde{B}^{(0)}=\alpha_{B}B_{\text{init}}\end{aligned}$ $\tilde{A}^{(0)}=\alpha_{A}A_{\text{init}},\tilde{B}^{(0)}=\alpha_{B}B_{\text{init}}\end{aligned}$
[…]
@@ 2.3 Magnitude Limitation Rooted in Low-Rank Structure
$\nu[W^{(t)}_{\text{LoRA}}]\approx k_{1}\gamma t+\mathcal{O}(\gamma^{2}t^{2}),\ \text{where}\ \gamma=\eta^{2}\sigma_{L},\ k_{1}=r(m\sigma_{A}^{4}+n\sigma_{B}^{4}).$ (6)
@@ 3.2 Efficient Magnitude-driven Initialization with LoRAM
$\Phi_{m}[i,j]=\sqrt{\frac{2}{m+1}}\sin\left(\frac{(i+1)(j+1)\pi}{m+1}\right),\quad 1\leq i,j\leq m.$
@@
$\Phi_{n}$, $\Phi_{m}\leftarrow$ get_basis(n, r), get_basis(m, r)
  // Generate basis matrices, Eq (12)
$\beta\leftarrow\left(\frac{\log_{\min(n,m)}(r)\cdot\nu[W]}{\nu[\Phi_{n}\Phi_{m}^{\top}]}\right)^{1/4}$
  // Compute magnitude gain factor
$B^{(0)},A^{(0)},W\leftarrow\beta\cdot\Phi_{n},\beta\cdot\Phi^{\top}_{m},W-\beta^{2}\cdot\Phi_{n}\Phi^{\top}_{m}$
  // Initialize parameters
@@ 4 Experiments
Baselines. While extensive research on LoRA has explored aspects like structural modifications and rank control, these directions are largely orthogonal to our focus on hyperparameter analysis within the naive LoRA framework. In line with this, we compare LoRAM with the naive LoRA (ICLR 2022) Hu2021LoRALA , as well as several representative hyperparameter tuning strategies. We first consider weight-driven initialization (marked “§$\S$”), including PiSSA (NeurIPS 2024) PiSSA , which uses the top-$r$ singular vectors and values of pre-trained weights; MiLoRA (NAACL 2025) MiLoRA , which utilizes the last $r$ singular vectors and values; and OLoRA olora , which applies orthogonal initialization via QR decomposition. All these methods adopt a fixed scaling factor $\alpha=1$. We then include RsLoRA kalajdzievski2023rank  (marked $\dagger$”), which enhances performance by setting $\alpha=\sqrt{r}$, and LoRA+ (ICML 2024) lora+  (marked $\ddagger$”), which increases the learning rate with the recommended $\eta_{B}=4\eta_{A}$. We also evaluate data-driven initialization in the ablation study, including LoRA-GA (NeurIPS 2024) wang2024lora and CorDA (NeurIPS 2024) corda , which require extra pipeline to leverage training data information. Most of these baselines have been integrated and validated in the PEFT library.
@@
Rank ##\##Param Method GSM8K MATH HumanEval MBPP Commonsense
@@
Method ##\##Param MNLI SST-2 MRPC CoLA QNLI QQP RTE STS-B
@@ Appendix A Related Work
LoRA (Hu2021LoRALA, ) has garnered significant attention in leveraging low-rank structures to represent weight updates, with subsequent studies expanding its foundations and applications (yang2024low, ; mao2025survey, ). Existing improvements to LoRA primarily focus on three directions:
@@
Proper initialization critically influences neural network training outcomes glorot2010understanding ; he2015delving . The standard LoRA (Hu2021LoRALA, ) initializes its low-rank matrices with random noise and zeros, referred to as "Noise & Zero" scheme, demonstrated to hinder convergence. Recent advances address this limitation by leveraging knowledge from pre-trained weights (PiSSA, ; MiLoRA, ; olora, ; SLoRA, ) or task-specific data (lora-one, ; corda, ; lora-ga, ). Most prominent approaches involve Singular Value Decomposition (SVD), which allows flexible control over matrix rank. As a pioneer, PiSSA PiSSA initializes LoRA weights using principal singular components of pre-trained matrices, aligning adaptation directions with the most significant parameter variations to accelerate convergence. Subsequent works corda ; lora-ga ; lora-one ; paischer2024one propose data-driven strategies that incorporate domain knowledge into adapter construction. For instance, LoRA-GA (lora-ga, ) aligns low-rank gradient directions with full fine-tuning counterparts during initialization. While these methods boost performance and retain the core LoRA structure, their theoretical effects on optimization dynamics remain unclear. More importantly, they are less pluggable than LoRA, requiring extra computational pipelines and storage for SVD buffers.
@@
Optimization Dynamics of LoRA. LoRA exhibits inherently nonlinear and non-convex optimization dynamics that complicate theoretical analysis (li2024crucial, ). Most existing theoretical studies are limited to simplified and idealized scenarios, such as the lazy-training regime jang2024lora ; malladi2023kernel and infinite-width limitlora+ ; hayou2024impact . A common goal across existing studies involves ensuring feature learning stability to prevent unstable or collapsed training dynamics. For instance,  kalajdzievski2023rank proves that improper scaling factors induce gradient collapse in high-rank adapters, proposing modified scaling mechanisms to stabilize forward and backward propagation. Recent work lora+ ; zhu2024asymmetry ; hayou2024impact reveals critical asymmetries in LoRA: The two low-rank matrices exhibit divergent distinct impacts on optimization trajectories. This asymmetry motivates lora+ to employ distinct learning rates for each matrix. For SVD-based initialization, lora-one investigates principal components derived from single-step full fine-tuning gradients lora-ga , providing convergence guarantees yet under restrictive assumptions. Moreover, such analyses still fail to explain the efficacy of alternative approaches like weight-driven initialization (PiSSA, ).
@@
The intricate nature of LoRA and its numerous advancements motivates us to seek a streamlined principle explaining its empirical success and guiding practical applications. We identify update magnitude amplification as a key mechanism, unifying seemingly disparate elements, such as scaling factors (kalajdzievski2023rank, ; biderman2024lora, ), initialization strategies (PiSSA, ; lora-ga, ), and learning rates (lora+, ), into a cohesive perspective.
@@ Appendix B Lower Bound on Representation Error
=$\mathcal{H}=\left\{(W+\alpha BA)x\mid B\in\mathbb{R}^{n\times r},A\in\mathbb{R}^{r\times m},\nu[A]\leq M_{1},\nu[B]\leq M_{2}\right\}.$
@@ C.1 Proof of Parameter Scaling Equivalence
Initialization $A^{(0)}=A_{\text{init}},B^{(0)}=B_{\text{init}}\end{aligned}$ $\tilde{A}^{(0)}=\alpha_{A}A_{\text{init}},\tilde{B}^{(0)}=\alpha_{B}B_{\text{init}}\end{aligned}$ $\tilde{A}^{(0)}=\alpha_{A}A_{\text{init}},\tilde{B}^{(0)}=\alpha_{B}B_{\text{init}}\end{aligned}$
[…]
@@ Appendix D Analyzing LoRA-GA from Magnitude Principle
We seek to determine the rank-$r$ matrices $A\in\mathbb{R}^{r\times m}$ and $B\in\mathbb{R}^{n\times r}$ that maximize the Frobenius norm of the gradients $\|\nabla_{A}L\|_{F}^{2}$ and $\|\nabla_{B}L\|_{F}^{2}$, subject to the constraints $\|A\|_{F}^{2}=r$ and $\|B\|_{F}^{2}=r$. From Eq. (2), the LoRA gradient update follows:
$\nabla_{A}L=\nabla_{W}LB^{\top},\quad\nabla_{B}L=\nabla_{W}^{\top}LA^{\top}.$ (56)
By substituting the singular value decomposition of $\nabla_{W}L$, given by
$\nabla_{W}L=USV^{\top},$ (57)
we obtain
$\nabla_{A}L=USV^{\top}B^{\top},\quad\nabla_{B}L=VSU^{\top}A^{\top}.$ (58)
Taking the Frobenius norm and using the cyclic property of the trace, we have
$\|\nabla_{A}L\|_{F}^{2}=\text{tr}(BS^{2}B^{\top}),\quad\|\nabla_{B}L\|_{F}^{2}=\text{tr}(AS^{2}A^{\top}).$ (59)
To maximize these expressions under the constraints $\|A\|_{F}^{2}=r$ and $\|B\|_{F}^{2}=r$, we use the variational characterization of the singular value decomposition. The optimal choice is to align $A$ and $B$ with the dominant singular vectors of $\nabla_{W}L$, leading to
$A^{\star}=V_{:,:r}^{\top},\quad B^{\star}=U_{:,:r}.$ (60)
This choice ensures that the gradient magnitude is maximized while satisfying the given constraints, completing the proof. ∎

### v2

@@ Abstract
Low-Rank Adaptation (LoRA) offers a parameter-efficient paradigm for tuning large models. While recent spectral initialization methods improve convergence and performance over the naive “Noise & Zeros” scheme, their extra computational and storage overhead undermines efficiency. In this paper, we establish update magnitude as the fundamental driver of LoRA performance and propose LoRAM, a magnitude-driven “Basis & Basis” initialization scheme that matches spectral methods without their inefficiencies [^1] Code is available here. Our key contributions are threefold: (i) Magnitude of weight updates determines convergence. We prove low-rank structures intrinsically bound update magnitudes, unifying hyperparameter tuning in learning rate, scaling factor, and initialization as mechanisms to optimize magnitude regulation. (ii) Spectral initialization succeeds via magnitude amplification. We demystify that the presumed knowledge-driven benefit of the spectral component essentially arises from the boost in the weight update magnitude. (iii) A novel and compact initialization strategy, LoRAM, scales deterministic orthogonal bases using pretrained weight magnitudes to simulate spectral gains. Extensive experiments show that LoRAM serves as a strong baseline, retaining the full efficiency of LoRA while matching or outperforming spectral initialization across benchmarks.
@@ 1 Introduction
The rise of large pretrained models Brown2020LanguageMA; OpenAI2023GPT4TR; touvron2023llama; bai2023qwen; guo2025deepseek has driven urgent needs for parameter-efficient fine-tuning (PEFT) methods peft; Hu2021LoRALA; Lester2021ThePO; He2021TowardsAU; Edalati2022KronAPE; zhang2025parameter. Among these, Low-Rank Adaptation (LoRA) Hu2021LoRALA stands out for its efficiency, flexibility, and stability. By freezing pretrained weights and injecting trainable low-rank matrices, LoRA enables the update of less than 1% of the parameters, significantly reducing memory and compute costs. Its plug-and-play nature achieves easy integration into diverse models, facilitating model sharing and federated learning SLoRA; FeDeRA. Additionally, LoRA helps prevent catastrophic forgetting biderman2024lora, making it well-suited for continual learning. These advantages have led to its wide adoption in multilingual NLP liu2022few; Ding2022DeltaTA; sun2023comparative; zhao2024lora and multimodal applications ye2023mplug; guo2023animatediff; blattmann2023stable; ruiz2023dreambooth.
@@ 2.1 Preliminaries and Notations
$y=Wx+W_{\text{LoRA}}x=Wx+\alpha(BA-B^{(0)}A^{(0)})x,$
where $B\in\mathbb{R}^{n\times r}$, $A\in\mathbb{R}^{r\times m}$ are trainable low-rank matrices with $r\ll\min(n,m)$, and $\alpha$ scales the update magnitude. The initialization term $B^{(0)}A^{(0)}$ can be absorbed into $W$ for the convenience, as illustrated in Figure 1(b). Given a loss function $L$, LoRA updates are computed as:
@@ 2.2 Effect of Hyperparameters on Update Magnitude
Initialization $\begin{aligned} A^{(0)}=A_{\text{init}},B^{(0)}=B_{\text{init}}\end{aligned}$ $\begin{aligned} \tilde{A}^{(0)}=\alpha_{A}A_{\text{init}},\tilde{B}^{(0)}=\alpha_{B}B_{\text{init}}\end{aligned}$ $\begin{aligned} \tilde{A}^{(0)}=\alpha_{A}A_{\text{init}},\tilde{B}^{(0)}=\alpha_{B}B_{\text{init}}\end{aligned}$
Learning Rates $\begin{aligned} \eta_{A}>0,\eta_{B}>0\end{aligned}$ $\begin{aligned} \eta_{\tilde{A}}=\alpha^{2}_{A}\eta_{A},\eta_{\tilde{B}}=\alpha^{2}_{B}\eta_{B}\end{aligned}$ $\begin{aligned} \eta_{\tilde{A}}=\alpha_{A}\eta_{A},\eta_{\tilde{B}}=\alpha_{B}\eta_{B}\end{aligned}$
@@ 2.3 Magnitude Limitation Rooted in Low-Rank Structure
$\nu[W^{(t)}_{\text{LoRA}}]\approx k_{1}\gamma t+\mathcal{O}(\gamma^{2}t^{2}),\ \text{where}\ \gamma=\eta^{2}\sigma_{L}^{2},\ k_{1}=r(m\sigma_{A}^{4}+n\sigma_{B}^{4}).$ (6)
@@ 3.2 Efficient Magnitude-driven Initialization with LoRAM
$\Phi_{m}[i,j]=\sqrt{\frac{2}{m+1}}\sin\left(\frac{(i+1)(j+1)\pi}{m+1}\right),\quad 0\leq i,j<m.$
@@
0.2cm $\Phi_{n}$, $\Phi_{m}\leftarrow$ get_basis(n, r), get_basis(m, r) $\triangleright$ Generate basis matrices, e.g., Eq (12)
$\beta\leftarrow\left(\frac{\log_{\min(n,m)}(r)\cdot\nu[W]}{\nu[\Phi_{n}\Phi_{m}^{\top}]}\right)^{1/4}$ $\triangleright$ Compute magnitude gain factor
$B^{(0)},A^{(0)},W\leftarrow\beta\cdot\Phi_{n},\beta\cdot\Phi^{\top}_{m},W-\beta^{2}\cdot\Phi_{n}\Phi^{\top}_{m}$ $\triangleright$ Initialize parameters
@@ 4 Experiments
Baselines. While extensive research on LoRA has explored aspects like structural modifications and rank control, these directions are largely orthogonal to our focus on hyperparameter analysis within the naive LoRA framework. In line with this, we compare LoRAM with the naive LoRA (ICLR 2022) Hu2021LoRALA, as well as several representative hyperparameter tuning strategies. We first consider weight-driven initialization (marked $\S$”), including PiSSA (NeurIPS 2024) PiSSA, which uses the top-$r$ singular vectors and values of pre-trained weights; MiLoRA (NAACL 2025) MiLoRA, which utilizes the last $r$ singular vectors and values; and OLoRA olora, which applies orthogonal initialization via QR decomposition. All these methods adopt a fixed scaling factor $\alpha=1$. We then include RsLoRA kalajdzievski2023rank (marked $\dagger$”), which enhances performance by setting $\alpha=\sqrt{r}$, and LoRA+ (ICML 2024) lora+ (marked $\ddagger$”), which increases the learning rate with the recommended $\eta_{B}=4\eta_{A}$. We also evaluate data-driven initialization in the ablation study, including LoRA-GA (NeurIPS 2024) wang2024lora and CorDA (NeurIPS 2024) corda, which require extra pipeline to leverage training data information. Most of these baselines have been integrated and validated in the PEFT library.
@@
Rank #\#Param Method GSM8K MATH HumanEval MBPP Commonsense
@@
Method #\#Param MNLI SST-2 MRPC CoLA QNLI QQP RTE STS-B
@@ Appendix A Related Work
The growing scale (Brown2020LanguageMA; OpenAI2023GPT4TR; touvron2023llama; bai2023qwen; karras2020analyzing; rombach2022high) of large pre-trained models across various domains urgently demands efficient fine-tuning methods (zhao2024lora; ye2023mplug; guo2023animatediff; blattmann2023stable; ruiz2023dreambooth; goga; ei2; li2025top; han2025joyagents). LoRA (Hu2021LoRALA) has garnered significant attention in leveraging low-rank structures to represent weight updates, with subsequent studies expanding its foundations and applications (yang2024low; mao2025survey). Existing improvements to LoRA primarily focus on three directions:
  • The low-rank constraint creates bottlenecks when learning complex features. Recent approaches enhance expressiveness through iterative stacking of LoRA modules (lialin2023relora; xia2024chain; Zi2023DeltaLoRAFH; PeriodicLoRA), high-order matrix operations (hyeon2021fedpara; Edalati2022KronAPE), and customized architectural designs (journals/corr/abs-2402-17263; jiang2024mora; li2024loran; zhong2024neat; Flora; Zhang2023AdaptiveBA; liu2024dora).
  • The non-convex landscapes pose challenges for numerical optimization and hyperparameter tuning (li2024crucial). Researchers have explored to improve scaling factor (kalajdzievski2023rank; Sebastian2023; biderman2024lora), learning rate (lora+), dropout rate (lora-dropout), optimizers (li2024crucial; zhang2024riemannian; lora-one; lora-ga), and initialization strategies (PiSSA; MiLoRA; olora; hayou2024impact; lora-one; corda; lora-ga).
  • Recent works reduce computational and memory overhead through freezing parameters (lora-fa; zhu2024asymmetry), designing more compact adapters (vera; lora-xs; gao2024parameter) and parameter quantization (PiSSA; QLoRA; LoftQ).
@@
Proper initialization critically influences neural network training outcomes glorot2010understanding; he2015delving. The standard LoRA (Hu2021LoRALA) initializes its low-rank matrices with random noise and zeros, referred to as "Noise & Zero" scheme, demonstrated to hinder convergence. Recent advances address this limitation by leveraging knowledge from pre-trained weights (PiSSA; MiLoRA; olora; SLoRA) or task-specific data (lora-one; corda; lora-ga). Most prominent approaches involve Singular Value Decomposition (SVD), which allows flexible control over matrix rank. As a pioneer, PiSSA PiSSA initializes LoRA weights using principal singular components of pre-trained matrices, aligning adaptation directions with the most significant parameter variations to accelerate convergence. Subsequent works corda; lora-ga; lora-one; paischer2024one propose data-driven strategies that incorporate domain knowledge into adapter construction. For instance, LoRA-GA (lora-ga) aligns low-rank gradient directions with full fine-tuning counterparts during initialization. While these methods boost performance and retain the core LoRA structure, their theoretical effects on optimization dynamics remain unclear. More importantly, they are less pluggable than LoRA, requiring extra computational pipelines and storage for SVD buffers.
@@
Optimization Dynamics of LoRA. LoRA exhibits inherently nonlinear and non-convex optimization dynamics that complicate theoretical analysis (li2024crucial). Most existing theoretical studies are limited to simplified and idealized scenarios, such as the lazy-training regime jang2024lora; malladi2023kernel and infinite-width limitlora+; hayou2024impact. A common goal across existing studies involves ensuring feature learning stability to prevent unstable or collapsed training dynamics. For instance,  kalajdzievski2023rank proves that improper scaling factors induce gradient collapse in high-rank adapters, proposing modified scaling mechanisms to stabilize forward and backward propagation. Recent work lora+; zhu2024asymmetry; hayou2024impact reveals critical asymmetries in LoRA: The two low-rank matrices exhibit divergent distinct impacts on optimization trajectories. This asymmetry motivates lora+ to employ distinct learning rates for each matrix. For SVD-based initialization, lora-one investigates principal components derived from single-step full fine-tuning gradients lora-ga, providing convergence guarantees yet under restrictive assumptions. Moreover, such analyses still fail to explain the efficacy of alternative approaches like weight-driven initialization (PiSSA).
@@
The intricate nature of LoRA and its numerous advancements motivates us to seek a streamlined principle explaining its empirical success and guiding practical applications. We identify update magnitude amplification as a key mechanism, unifying seemingly disparate elements, such as scaling factors (kalajdzievski2023rank; biderman2024lora), initialization strategies (PiSSA; lora-ga), and learning rates (lora+), into a cohesive perspective.
@@ Appendix B Lower Bound on Representation Error
$\mathcal{H}=\left\{(W+\alpha BA)x\mid B\in\mathbb{R}^{n\times r},A\in\mathbb{R}^{r\times m},\nu[A]\leq M_{1},\nu[B]\leq M_{2}\right\}.$
@@ C.1 Proof of Parameter Scaling Equivalence
Initialization $\begin{aligned} A^{(0)}=A_{\text{init}},B^{(0)}=B_{\text{init}}\end{aligned}$ $\begin{aligned} \tilde{A}^{(0)}=\alpha_{A}A_{\text{init}},\tilde{B}^{(0)}=\alpha_{B}B_{\text{init}}\end{aligned}$ $\begin{aligned} \tilde{A}^{(0)}=\alpha_{A}A_{\text{init}},\tilde{B}^{(0)}=\alpha_{B}B_{\text{init}}\end{aligned}$
Learning Rates $\begin{aligned} \eta_{A}>0,\eta_{B}>0\end{aligned}$ $\begin{aligned} \eta_{\tilde{A}}=\alpha^{2}_{A}\eta_{A},\eta_{\tilde{B}}=\alpha^{2}_{B}\eta_{B}\end{aligned}$ $\begin{aligned} \eta_{\tilde{A}}=\alpha_{A}\eta_{A},\eta_{\tilde{B}}=\alpha_{B}\eta_{B}\end{aligned}$
@@ Appendix D Analyzing LoRA-GA from Magnitude Principle
Let $G=\nabla_{W}L\in\mathbb{R}^{n\times m}$ and write its SVD $G=USV^{\top}$ with singular values $\sigma_{1}\geq\sigma_{2}\geq\cdots\geq 0$. From the LoRA gradient expressions we have
$\nabla_{A}L=B^{\top}G,\qquad\nabla_{B}L=GA^{\top}.$
Note that $\nabla_{A}L$ does not depend on $A$ and $\nabla_{B}L$ does not depend on $B$, hence the two maximization problems decouple.
(i) Maximizing $\|\nabla_{B}L\|_{F}^{2}$ w.r.t. $A$ under $\|A\|_{F}^{2}=r$. Let $A^{\prime}:=AV$. Then $\|A^{\prime}\|_{F}^{2}=\|A\|_{F}^{2}=r$ and
$\|GA^{\top}\|_{F}^{2}=\|US(A^{\prime})^{\top}\|_{F}^{2}=\|S(A^{\prime})^{\top}\|_{F}^{2}=\mathrm{tr}\big(A^{\prime}S^{2}(A^{\prime})^{\top}\big).$
Form the Lagrangian $\mathcal{L}(A^{\prime},\lambda)=\mathrm{tr}(A^{\prime}S^{2}(A^{\prime})^{\top})-\lambda(\mathrm{tr}(A^{\prime}(A^{\prime})^{\top})-r)$. Taking derivative w.r.t. $A^{\prime}$ yields the stationarity condition $A^{\prime}S^{2}=\lambda A^{\prime}$. Thus the rows of $A^{\prime}$ lie in the eigenspaces of $S^{2}$. To maximize the trace, one places the Frobenius norm onto the coordinates corresponding to the largest diagonal entries of $S^{2}$, i.e. the first $r$ singular values. Taking $A^{\prime\star}=[I_{r}\mid 0]$ attains the maximum
$\max_{\|A\|_{F}^{2}=r}\|GA^{\top}\|_{F}^{2}=\sum_{i=1}^{r}\sigma_{i}^{2},$
and $A^{\star}=A^{\prime\star}V^{\top}=V_{:,:r}^{\top}$ (up to orthonormal transformations within the chosen $r$-subspace).
(ii) Maximizing $\|\nabla_{A}L\|_{F}^{2}$ w.r.t. $B$ under $\|B\|_{F}^{2}=r$. Analogously let $B^{\prime}:=U^{\top}B$. Then
$\|B^{\top}G\|_{F}^{2}=\mathrm{tr}\big((B^{\prime})^{\top}S^{2}B^{\prime}\big),$
and the same Lagrange argument gives that the optimal choice concentrates norm on the first $r$ coordinates, producing
$\max_{\|B\|_{F}^{2}=r}\|B^{\top}G\|_{F}^{2}=\sum_{i=1}^{r}\sigma_{i}^{2},$
and one may take $B^{\star}=U_{:,:r}$.
Hence the stated solutions $A^{\star}=V_{:,:r}^{\top}$ and $B^{\star}=U_{:,:r}$ are optimal, with the above maximal values. This completes the proof. ∎